Operation Summary for December 14, 2009 - December 16, 2009
Monday Dec. 14
During the day shift, at 9:50 AM, NuMI personnel accessed their Absorber Hall to investigate a problem with an expansion tank skid. Operations terminated store 7437 at 9:55 AM and then put the TeV into HEP shot setup. NuMI personnel came out of access at 10:50 AM. Operations established store 7438 at 11:15 AM with an average initial luminosity (AIL) of 255.0E30.
During the evening shift, Operations reported that a Recycler kicker power supply (KPS2A) tripped off; EE Support personnel lowered the reservoir voltage to return the kicker to service. Operations disabled the MTest beam at 9 PM.
Tuesday Dec. 15
During the midnight shift, the Booster East Anode Power Supply (EAPS) tripped off at 3:15 AM due to a transformer problem on a RF station (BRF2). Operators bypassed the station and brought BRF19 online. Operations called in Booster RF personnel to work on BRF2 and 8; they brought BRF8 online. Operations terminated store 7438 at 4:20 AM and then put the TeV into shot setup. Operations established store 7439 at 5:50 AM with an AIL of 249.0E30. Operations established beam to MTest at 6:20 AM.
During the day shift, TeV personnel conducted beam studies during the morning. Operations turned off a TeV RF station (TRF5) due to output gap envelope problems.
During the evening shift, TeV personnel began an end-of-store study at 7 PM. Operations terminated store 7439 at 10 PM. Operations brought TRF5 back online. Operations put the TeV into shot setup at 10:15 PM and established store 7440 at 11 PM with an AIL of 227.0E30. Operations disabled the MTest beam at 6 PM.
Wednesday Dec. 16
During the midnight shift, at 12:25 AM, the Booster EAPS tripped off due to a BRF8 modulator problem. A Booster vertical trim magnet (B:VS4) current drifted, causing beam position problems. Operations restored the settings to the power supply’s control crate, which corrected the beam position. At 4:05 AM, the Booster EAPS tripped off due to a BRF1 water pressure sensor. Booster RF personnel came in and returned the station to service.
The Plans for Wednesday
Comments and Suggestions
|last modified 12/16/2009 email Fermilab|
|Security, Privacy, Legal|